# Singing Voice Synthesis Using Differentiable LPC and Glottal-Flow-Inspired Wavetables

### v1

@@ Abstract
This paper introduces GlOttal-flow LPC Filter (GOLF), a novel method for singing voice synthesis (SVS) that exploits the physical characteristics of the human voice using differentiable digital signal processing. GOLF employs a glottal model as the harmonic source and IIR filters to simulate the vocal tract, resulting in an interpretable and efficient approach. We show it is competitive with state-of-the-art singing voice vocoders, requiring fewer synthesis parameters and less memory to train, and runs an order of magnitude faster for inference. Additionally, we demonstrate that GOLF can model the phase components of the human voice, which has immense potential for rendering and analysing singing voices in a differentiable manner. Our results highlight the effectiveness of incorporating the physical properties of the human voice mechanism into SVS and underscore the advantages of signal-processing-based approaches, which offer greater interpretability and efficiency in synthesis.
@@ 1 Introduction
In contrast, Differentiable Digital Signal Processing (DDSP) models [14, 15, 16] incorporate existing signal processing operations into neural networks as an inductive bias, making them more interpretable and generalisable. DDSP additive synthesis has been tried for SVS by Alonso et al. [17], and Wu et al. [18] improved it further by using subtractive synthesis and sawtooth as the harmonic source. Nercessian et al. [19] proposed a differentiable version of the WORLD vocoder [20] for doing end-to-end singing voice conversion. Yoshimura et al. [21] used Taylor expansion to approximate the mel-log spectrum approximation filter’s (MLSA) exponential function and embedded it into an SVS system. However, most of their architectures only assume the target signal is a monophonic instrument. The harmonic sources are fixed to a specific shape (e.g. sawtooth, pulse train), and the filters are symmetric in the time domain, except Yoshimura et al. [21] which use a minimum-phase MLSA filter. Incorporating constraints specific to the human voice on the harmonic source and the filters could lead to a more interpretable and compact SVS vocoder.
In this work, we propose GlOttal-flow LPC Filter (GOLF), an SVS module informed by the physical properties of the human voice. We build upon the Harmonic-plus-Noise architecture of DDSP [14] and the subtractive synthesis of SawSing [18] but replace the harmonic source with a glottal model and use IIR filters. We developed a differentiable IIR implementation in PyTorch [22] for training efficiency. We then used this module as a neural vocoder and compared its performance with other DDSP-based vocoders. Specifically, a simple and lightweight NN encoder converts the mel-spectrogram into synthesis parameters, and the synthesiser decodes the signal from it. We paired different synthesisers with the same encoder and trained them jointly.
Our contributions are twofold. First, GOLF has significantly fewer synthesis parameters but is still competitive with state-of-the-art SVS vocoders. Second, GOLF requires less than 40% of memory to train and runs ten times faster than its alternatives for inference. Moreover, we indirectly show that GOLF could model the phase components of the human voice by aligning the synthesised waveforms to the ground truth and calculating the differences. This characteristic has excellent potential for analysing singing voices in a differentiable manner. Decomposing the human voice into the glottal source and vocal tract could also enable us to adjust the singing style in different ways, such as altering the amount of vocal effort with varying shapes of the glottal pulse.
@@ 2.1 Glottal Source-Filter Model
where $G(z)$ represents the periodic vibration from the vocal folds, $N(z)$ represents random components of the glottal source, $H(z)$ represents the vocal-tract filter, and $L(z)$ represents the radiation at the lips [23]. Since this formulation is linear, the radiation filter $L(z)$ and the glottal pulse $G(z)$ can be merged into a single source $G^{\prime}(z)$ called the radiated glottal pulse. If we assume $L(z)$ is a first-order differentiator $1+z^{-1}$ [24], then $G^{\prime}(z)$ is the derivative of the glottal pulse, which can be described by the LF model [25], a four-parameter model of glottal flow. $H(z)$ is usually a Linear Predictive Coding (LPC) filter.
@@ 2.2 Linear Predictive Coding
where $a_{i}$ are the linear prediction coefficients. This is the same as filtering the residuals, equivalent to the glottal source in our case, with an $K^{th}$ order all-pole filter, a filter that has an infinite impulse response (IIR). We can use the LPC filter to represent the response of the vocal tract if the vocal tract is approximated by a series of cylindrical tubes with varying diameters [26], providing a physical interpretation.
@@ 3 Proposed Model
Usually, $N(z)$ in Eqn (1) is treated as amplitude-modulated Gaussian noise [24, 23]. Our early experiments found this formulation to be challenging to optimise. As an alternative, we move the noise components $N(z)$ outside the glottal source and filter it with time-varying $C(z)$, resulting in
@@
This resembles the classic Harmonic-plus-Noise model [36] and was used in previous DDSP-based SVS [17, 18]. Alonso et al. [17] modelled $G^{\prime}(z)H(z)$ jointly using additive harmonic oscillators and time-varying Finite Impulse Responses (FIRs) as $C(z)$; Wu et al. [18] further introduced a sawtooth oscillator as $G^{\prime}(z)$ and zero-phase time-varying FIRs as $H(z)$. In this work, we use a glottal flow model to synthesise harmonic sources and time-varying IIRs for filters.
@@ 3.1 Glottal Flow Wavetables
We adopted the transformed-LF model [37] for generating glottal pulses. This model re-parameterises the LF model [25] using just one parameter $R_{d}$, which has been found to correspond well to the perceived vocal effort and covers a wide range of different glottal flow shapes. We sampled $K$ value of $\log(R_{d})$ with equal spacing inside $[\log(0.3),\log(2.7)]$ according to the value range suggested by [23]. We calculate the flow derivative function $g^{\prime}(t;R_{d})$ in continuous time $t$ for each sampled $R_{d}$ and then sampled $L$ points in one period to get its discrete version. The details for calculating $g^{\prime}(t;R_{d})$ were given in [38]. By stacking these sampled glottal flows, we built wavetables $\mathbf{D}\in\mathbb{R}^{K\times L}$, with each row containing one period of a sampled glottal pulse (see Fig. 1). The rows are sorted using $R_{d}$.
@@ 3.2 Frame-Wise LPC Synthesis
Time-varying LPC synthesis is usually done by linearly interpolating the LPC coefficients to the audio resolution and filtering sample by sample. This is not parallelisable and slows down the training process. As an alternative, we approximate LPC synthesis by doing it on the frame level and using overlap-add:
@@
where $LPC(e_{n};\mathbf{a})$ means Eqn (2), ${\mathbf{a}}_{k}\in\mathbb{R}^{M}$ are the filter coefficients at the $k^{th}$ frame, $u_{n}$ and $w_{n}$ are the windowing functions, $\gamma_{n}\in\mathbb{R}^{+}$ is the gain, and $T$ is the hop size. $u_{n}$ is fixed to the square window. In this way, every frame is independent of others and can be parallelised. We found that the voice quality differences between overlap-add LPC and sample-by-sample LPC are barely noticeable if we use a sufficiently small hop size. We found 200 Hz frame rate is enough, empirically.
@@ 3.3 LPC Coefficients Parameterisation
To make the LPC filter stable, all of its poles must lie inside the unit circle on the complex plane. The stability can be guaranteed using robust representations, such as reflection coefficients [28]. The representation we chose in this work is cascaded 2nd-order IIR filters, and we solve the stability issue by ensuring all the 2nd-order filters are stable. For an $M^{th}$ order LPC filter, it needs $\frac{M}{2}$ such filters. We use the coefficient representation from [33] to parameterise the $i^{th}$ IIR filter’s coefficients $1+\eta_{i,1}z^{-1}+\eta_{i,2}z^{-2}$ from the encoder’s outputs and cascade them together to form the LPC coefficients:
@@ 3.4 Unvoiced Gating
The instantaneous frequency $f_{n}$ predicted by the encoder is always non-zero and keeps the oscillator working. Without constraint, the model would utilise these harmonics in the unvoiced region, creating buzzing artefacts [18]. We proposed to mitigate this problem by jointly training the model to predict the voiced regions with $v_{n}\in[0,1]$ and feeding the gated frequency $\hat{f}_{n}=v_{n}f_{n}$ to the oscillator instead.
@@ 4 Optimisation
Training deep learning models is usually done by backpropagating the gradients evaluated at a chosen loss function $\mathcal{L}$ throughout the whole computational graph back to the parameters. Partially inspired by Bhattacharya et al.[29], we derived the closed formation of backpropagation through time to utilise efficient IIR implementation to solve the problems we mentioned in Section 2.2 while keeping the filter differentiable. Here, $\mathbf{e}\in\mathbb{R}^{N}$ is the input, ${\mathbf{a}}\in\mathbb{R}^{M}$ is the filter coefficients, and $\mathbf{s}\in\mathbb{R}^{N}$ is the output. Assuming we know $\frac{\partial\mathcal{L}}{\partial\mathbf{s}}$, we can get the derivatives $\frac{\partial\mathcal{L}}{\partial\mathbf{e}}$ and $\frac{\partial\mathcal{L}}{\partial{\mathbf{a}}}$, using chain rules $\frac{\partial\mathcal{L}}{\partial\mathbf{s}}\frac{\partial\mathbf{s}}{\partial\mathbf{e}}$ and $\frac{\partial\mathcal{L}}{\partial\mathbf{s}}\frac{\partial\mathbf{s}}{\partial{\mathbf{a}}}$.
@@ 4.1 Backpropagation Through the Coefficients
which equals $LPC(-s_{n-i};\mathbf{a})$. The initial conditions $\frac{\partial s_{n}}{\partial a_{i}}|_{n\leq 0}$ are zeros if the initial rest condition $s_{n}|_{n\leq 0}=0$ is used; for non-zero initial conditions, the derivatives can be derived in the same way with minor modifications. We can get $\frac{\partial s_{n}}{\partial{\mathbf{a}}}$ with one pass of filtering because $\frac{\partial s_{n}}{\partial a_{j}}$ is $\frac{\partial s_{n}}{\partial a_{i}}$ with a time-shift $j-i$.
@@ 4.2 Backpropagation Through the Input
In conclusion, backpropagation through an IIR filter consists of two passes of the same IIR and one matrix multiplication.  [^1] The computation of the IIR does not need to fulfil the implementation requirements set by the automatic differentiation framework, thus can be highly optimised. We implemented the IIR in C++ and CUDA with multi-threading to filter multiple sequences simultaneously.  [^2] Although the single-core performance of a GPU is usually inferior to a CPU, and we can only use at most one thread for each IIR, the GPU has a much higher number of cores, which is beneficial for training on a large number of sequences at once. The differentiable IIR is done by registering the above backward computation in PyTorch, and we submitted the implementation to Torchaudio [39] as part of the torchaudio.functional.lfilter.
@@ 5.1 Dataset
We test GOLF as a neural vocoder on the MPop600 dataset [40], a high-quality Mandarin singing voice dataset featuring nearly 600 singing recordings with aligned lyrics sung by four singers. We used the audio recordings from the f1 (female) and m1 (male) singers. For each singer, we selected the first three recordings as the test set, the following 27 recordings as the validation set, and used the rest as training data (around three hours in total). All the recordings were downsampled to 24 kHz. The vocoder feature we choose is the log mel-spectrogram. We computed the feature with a window size of 1024 and 80 mel frequency bins and set the hop size $T$ to 120. We normalised the feature to between zero and one and sliced the training data into two seconds excerpts with 1.5 seconds overlap.
@@ 5.2 Model Details
We adopted the encoder from SawSing but replaced the transformer layers with three layers of Bi-LSTM in favour of less implementation difficulty, resulting in around 0.7M parameters in total. A linear layer at the end predicts the synthesis parameters $\{f_{k},v_{k},\gamma_{k},\beta_{k},{\mathbf{a}}_{k},\mathbf{b}_{k}\}$. The first four parameters are linearly upsampled to $\{f_{n},v_{n},\gamma_{n},\beta_{n}\}$. $\beta_{n}$ and $\mathbf{b}_{k}$ are the Gaussian noise’s gain and filter coefficients. We added an average pooling layer with a size of 10 and two convolution layers after the encoder to predict the $R_{d}$ fractional index $\tau_{o}$ at a lower rate and then linearly upsampled to $\tau_{n}$. This step avoids possible modulation effects caused by switching the wavetables too quickly. A system diagram of GOLF is shown in Fig. 2. We set $K=100,L=2048$, Hanning window for $w_{n}$, and $M=22$ for both LPC filters. We used the same hop size $T$ and a window size of 480 for frame-wise LPC. We normalised all wavetables to have equal energy and aligned them with the negative peak.
@@
We compare GOLF with three DDSP-based baselines using the same NN encoder to predict their synthesis parameters. The first two are the original DDSP [14] and SawSing [18]. We set their noise filter length to 80 and harmonic filter length to 256 for SawSing. The third model is PUlse-train LPC Filter (PULF), which is GOLF but replaces the glottal flow wavetables with a band-limited pulse train [19] using additive synthesis, and the LPC order for the harmonic source is increased to 26 to accommodate the glottal pulse response. The number of oscillating sinusoids was set to over 150 for all the baselines. The reason we did not compare GOLF with Nercessian et al. [19] and Yoshimura et al. [21] is that their architectures are based closely on the source-filter model, and they used additional post-nets to enhance the voice, which makes it harder to compare directly with GOLF.

### v3

@@ Abstract
This paper introduces GlOttal-flow LPC Filter (GOLF), a novel method for singing voice synthesis (SVS) that exploits the physical characteristics of the human voice using differentiable digital signal processing. GOLF employs a glottal model as the harmonic source and IIR filters to simulate the vocal tract, resulting in an interpretable and efficient approach. We show it is competitive with state-of-the-art singing voice vocoders, requiring fewer synthesis parameters and less memory to train, and runs an order of magnitude faster for inference. Additionally, we demonstrate that GOLF can model the phase components of the human voice, which has immense potential for rendering and analysing singing voice in a differentiable manner. Our results highlight the effectiveness of incorporating the physical properties of the human voice mechanism into SVS and underscore the advantages of signal-processing-based approaches, which offer greater interpretability and efficiency in synthesis.
@@ 1 Introduction
In contrast, Differentiable Digital Signal Processing (DDSP) models [14, 15, 16] incorporate existing signal processing operations into neural networks as an inductive bias, making them more interpretable and generalisable. DDSP additive synthesis has been proposed for SVS by Alonso et al. [17]. Wu et al. [18] improved this further by using subtractive synthesis and sawtooth as the harmonic source. Nercessian et al. [19] proposed a differentiable version of the WORLD vocoder [20] for doing end-to-end singing voice conversion. Yoshimura et al. [21] used Taylor expansion to approximate the mel-log spectrum approximation filter’s (MLSA) exponential function and embedded it into an SVS system. However, most of their architectures only assume the target signal is a monophonic instrument, which can potentially lead to solutions that do not reflect some properties of voice. In their design, the harmonic sources are fixed to a specific shape (e.g. sawtooth, pulse train), and the filters are symmetric in the time domain, except Yoshimura et al. [21] which use a minimum-phase MLSA filter. Incorporating constraints specific to the human voice on the harmonic source and the filters could lead to a more interpretable and compact SVS vocoder.
In this work, we propose GlOttal-flow LPC Filter (GOLF), an SVS module informed by the physical properties of the human voice. We build upon the Harmonic-plus-Noise architecture of DDSP [14] and the subtractive synthesis of SawSing [18], but replace the harmonic source with a glottal model and use IIR filters. We developed a differentiable IIR implementation in PyTorch [22] for training efficiency. We then used this module as a neural vocoder and compared its performance with other DDSP-based vocoders. Specifically, a simple and lightweight NN encoder converts the mel-spectrogram into synthesis parameters, and the synthesiser decodes the signal from it. We paired different synthesisers with the same encoder and trained them jointly.
Our contributions are twofold. First, GOLF has significantly fewer synthesis parameters but is still competitive with state-of-the-art SVS vocoders. Second, GOLF requires less than 40% of memory to train and runs ten times faster than its alternatives for inference. Moreover, we indirectly show that GOLF could model the phase components of the human voice by aligning the synthesised waveforms to the ground truth and calculating the differences. This characteristic has excellent potential for analysing singing voice in a differentiable manner. Decomposing the human voice into the glottal source and vocal tract could also enable us to adjust the singing style in different ways, such as altering the amount of vocal effort with varying shapes of the glottal pulse.
@@ 2.1 Glottal Source-Filter Model
where $G(z)$ represents the periodic vibration from the vocal folds, $N(z)$ represents random components of the glottal source, $H(z)$ represents the vocal-tract filter, and $L(z)$ represents the radiation at the lips [23]. Since this formulation is linear, the radiation filter $L(z)$ and the glottal pulse $G(z)$ can be merged into a single source $G^{\prime}(z)$ called the radiated glottal pulse. If we assume $L(z)$ is a first-order differentiator $1-z^{-1}$ [24], then $G^{\prime}(z)$ is the derivative of the glottal pulse, which can be described by the LF model [25], a four-parameter model of glottal flow. $H(z)$ is usually a Linear Predictive Coding (LPC) filter.
@@ 2.2 Linear Predictive Coding
where $a_{i}$ are the linear prediction coefficients. This is the same as filtering the residuals, equivalent to the glottal source in our case, with an $M^{th}$-order all-pole filter, a filter that has an infinite impulse response (IIR). We can use the LPC filter to represent the response of the vocal tract if the vocal tract is approximated by a series of cylindrical tubes with varying diameters [26], providing a physical interpretation.
@@ 3 Proposed Model
Usually, $N(z)$ in Eqn (1) is treated as amplitude-modulated Gaussian noise [24, 23]. Our early experiments found this formulation to be challenging to optimise. As an alternative, we move the noise components $N(z)$ outside the glottal source and filter it with time-varying filter $C(z)$, resulting in
@@
This resembles the classic Harmonic-plus-Noise model [36] and was used in previous DDSP-based SVS [17, 18]. Alonso et al. [17] modelled $G^{\prime}(z)H(z)$ jointly using additive harmonic oscillators and time-varying Finite Impulse Responses (FIRs) as $C(z)$; Wu et al. [18] introduced a sawtooth oscillator as $G^{\prime}(z)$ and zero-phase time-varying FIRs as $H(z)$. In this work, we use a glottal flow model to synthesise harmonic sources and time-varying IIRs as filters.
@@ 3.1 Glottal Flow Wavetables
We adopted the transformed-LF model [37] for generating glottal pulses. This model re-parameterises the LF model [25] using just one parameter $R_{d}$, which has been found to correspond to the perceived vocal effort well and covers a wide range of different glottal flow shapes. We sampled $K$ values of $\log(R_{d})$ with equal spacing inside $[\log(0.3),\log(2.7)]$ according to the value range suggested by [23]. We calculate the flow derivative function $g^{\prime}(t;R_{d})$ in continuous time $t$ for each sampled $R_{d}$ and then sampled $L$ points in one period to get its discrete version. The details for calculating $g^{\prime}(t;R_{d})$ were given in [38]. By stacking these sampled glottal flows, we built wavetables $\mathbf{D}\in\mathbb{R}^{K\times L}$, with each row containing one period of a sampled glottal pulse (see Fig. 1). The rows are sorted based on $R_{d}$.
@@ 3.2 Frame-Wise LPC Synthesis
Time-varying LPC synthesis is usually done by linearly interpolating the LPC coefficients to the audio resolution and filtering sample by sample. This is not parallelisable and slows down the training process. As an alternative, we approximate LPC synthesis by treating each frame independently and using overlap-add:
@@
where $LPC(e_{n};\mathbf{a})$ represents Eqn (2), ${\mathbf{a}}_{k}\in\mathbb{R}^{M}$ are the filter coefficients at the $k^{th}$ frame, $u_{n}$ and $w_{n}$ are the windowing functions, $\gamma_{n}\in\mathbb{R}^{+}$ is the gain, and $T$ is the hop size. $u_{n}$ is fixed to the square window. In this way, the computation can be parallelised. We found that the voice quality differences between overlap-add LPC and sample-by-sample LPC are barely noticeable if we use a sufficiently small hop size. We empirically found that a 200 Hz frame rate is sufficient.
@@ 3.3 LPC Coefficients Parameterisation
For the LPC filter to be stable, all of its poles must lie inside the unit circle on the complex plane. Stability can be guaranteed using robust representations, such as reflection coefficients [28]. The representation we chose in this work is cascaded 2nd-order IIR filters, and we solve the stability issue by ensuring all the 2nd-order filters are stable. We use the coefficient representation from [33] to parameterise the $i^{th}$ IIR filter’s coefficients $1+\eta_{i,1}z^{-1}+\eta_{i,2}z^{-2}$ from the encoder’s outputs and cascade them together to form an $M^{th}$-order LPC filter:
@@ 3.4 Unvoiced Gating
The instantaneous frequency $f_{n}$ predicted by the encoder is always non-zero and keeps the oscillator working. Without constraint, the model would utilise these harmonics in the unvoiced region creating buzzing artefacts [18]. We propose to mitigate this problem by jointly training the model to predict the voiced/unvoiced probabilities as $v_{n}\in[0,1]$ and feeding the gated frequency $\hat{f}_{n}=v_{n}f_{n}$ to the oscillator instead.
@@ 4 Optimisation
Training deep learning models is usually accomplished by backpropagating the gradients evaluated at a chosen loss function $\mathcal{L}$ throughout the whole computational graph back to the parameters. Partially inspired by Bhattacharya et al.[29], we derived the closed form of backpropagation through time to utilise efficient IIR implementation to solve the problems we mentioned in Section 2.2 while keeping the filter differentiable. Here, $\mathbf{e}\in\mathbb{R}^{N}$ is the input, ${\mathbf{a}}\in\mathbb{R}^{M}$ is the filter coefficients, and $\mathbf{s}\in\mathbb{R}^{N}$ is the output. Assuming we know $\frac{\partial\mathcal{L}}{\partial\mathbf{s}}$, we can get the derivatives $\frac{\partial\mathcal{L}}{\partial\mathbf{e}}$ and $\frac{\partial\mathcal{L}}{\partial{\mathbf{a}}}$, using chain rules $\frac{\partial\mathcal{L}}{\partial\mathbf{s}}\frac{\partial\mathbf{s}}{\partial\mathbf{e}}$ and $\frac{\partial\mathcal{L}}{\partial\mathbf{s}}\frac{\partial\mathbf{s}}{\partial{\mathbf{a}}}$.
@@ 4.1 Backpropagation Through the Coefficients
which equals $LPC(-s_{n-i};\mathbf{a})$. $s_{n}|_{n\leq 0}$ does not depend on $a_{i}$ so the initial conditions $\frac{\partial s_{n}}{\partial a_{i}}|_{n\leq 0}$ are zeros. We can get $\frac{\partial s_{n}}{\partial{\mathbf{a}}}$ with one pass of filtering because $\frac{\partial s_{n}}{\partial a_{j}}$ is $\frac{\partial s_{n}}{\partial a_{i}}$ shifted by an offset $j-i$. Lastly, we calculate $\frac{\partial\mathcal{L}}{\partial a_{i}}$ as $\sum_{n=1}^{N}\frac{\partial\mathcal{L}}{\partial s_{n}}\frac{\partial s_{n}}{\partial a_{i}}$.
@@ 4.2 Backpropagation Through the Input
In conclusion, backpropagation through an IIR filter consists of two passes of the same filter and one matrix multiplication [^1] The computation of the IIR does not need to fulfil the implementation requirements set by the automatic differentiation framework, thus can be highly optimised.. We implemented the IIR in C++ and CUDA with multi-threading to filter multiple sequences simultaneously [^2] Although the single-core performance of a GPU is usually inferior to a CPU, and we can only use at most one thread for each IIR, the GPU has a much higher number of cores, which is beneficial for training on a large number of sequences at once.. The differentiable IIR is done by registering the above backward computation in PyTorch, and we submitted the implementation to TorchAudio [39] as part of the torchaudio.functional.lfilter.
@@ 5.1 Dataset
We test GOLF as a neural vocoder on the MPop600 dataset [40], a high-quality Mandarin singing voice dataset featuring nearly 600 singing recordings with aligned lyrics sung by four singers. We used the audio recordings from the f1 (female) and m1 (male) singers. For each singer, we selected the first three recordings as the test set, the following 27 recordings as the validation set, and used the rest as training data (around three hours in total). All the recordings were downsampled to 24 kHz. The vocoder feature we choose is the log mel-spectrogram. We computed the feature with a window size of 1024 and 80 mel-frequency bins and set the hop size $T$ to 120. We normalised the feature to between zero and one and sliced the training data into two seconds excerpts with 1.5 seconds overlap.
@@ 5.2 Model Details
We adopted the encoder from SawSing but replaced the transformer layers with three layers of Bi-LSTM for favourable implementation, resulting in around 0.7M parameters in total. A final linear layer predicts the synthesis parameters $\{f_{k},v_{k},\gamma_{k},\beta_{k},{\mathbf{a}}_{k},\mathbf{b}_{k}\}$. The first four parameters are linearly upsampled to $\{f_{n},v_{n},\gamma_{n},\beta_{n}\}$. $\beta_{n}$ and $\mathbf{b}_{k}$ are the Gaussian noise’s gain and filter coefficients. We added an average pooling layer with a size of 10 and two convolution layers after the encoder to predict the $R_{d}$ fractional index $\tau_{o}$ at a lower rate and then linearly upsampled to $\tau_{n}$. This step avoids possible modulation effects caused by switching the wavetables too quickly. A system diagram of GOLF is shown in Fig. 2. We set $K=100,L=2048$, Hanning window for $w_{n}$, and $M=22$ for both LPC filters. We used the same hop size $T$ and a window size of 480 for frame-wise LPC. We normalised all wavetables to have equal energy and aligned them with the negative peak.
@@
We compare GOLF with three DDSP-based baselines using the same NN encoder to predict their synthesis parameters. The first two are the original DDSP [14] and SawSing [18]. We set their noise filter length to 80 and harmonic filter length to 256 for SawSing. The third model is PUlse-train LPC Filter (PULF), which is similar to GOLF but replaces the glottal flow wavetables with a band-limited pulse train [19] using additive synthesis, while the LPC order for the harmonic source is increased to 26 to accommodate the glottal pulse response. The number of oscillating sinusoids was set to over 150 for all the baselines. We did not compare GOLF with Nercessian et al. [19] and Yoshimura et al. [21] because these architectures are based closely on the source-filter model, and use additional post-nets to enhance the voice, which makes it harder to compare directly with GOLF.